Solve:
sqrt(x+1) = sqrt(2x-6)


-sqrt(y) - 8 = 6
: Solve:
sqrt(x+1) = sqrt(2x-6)


-sqrt(y) - 8 = 6

Answer by Alan3354(1214)

sqrt(x+1) = sqrt(2x-6)
-----------------------
Square both sides
x+1 = 2x-6
1 = x-6
7 = x
x = 7
-----------
-sqrt(y) - 8 = 6
-sqrt(y) = 14
y = 196